OpenStudy (anonymous):

In the 30-60-90 triangle below, side s has a length of _____ and side r has a length of _____.

OpenStudy (anonymous):



OpenStudy (anonymous):

so we need to find x we know that the hypot is 2x so if we set that equal to each other we get 2x = 12 x = 12/2 = 6 now we have 6 = x s = x and\[r = x \sqrt{3}\] here is a picture to help you understand what i just said